The total revenue of Retailers,Inc. increased according to the following table:
Year x since 1997 Total Revenue, R (millions $)
1997, x = 0 $22

1998, x = 1 $23

1999, x = 2 $27

2000, x = 3 $31

2001, x = 4 $33

2002, x = 5 $38

Use the points (1,23) and (5,38) to find the equation of the function that fits this data.
Now estimate the total revenue in the year 2008.

Answer by stanbon(75887)

2002, x = 5 $38
Use the points (1,23) and (5,38) to find the equation of the function that fits this data.
Now estimate the total revenue in the year 2008.
----------------
slope = [38-23]/[5-1] = 15/4
If y=23 when x=1 and slope = 15/4, then:
23 = (15/4)1+b
b= (92-15)/4
b= 77/4
EQUATION:
Revenue = (15/4)(year) + 77/4
"year" for 2008 = 2002+6 = 5+6=11
------------
Revenue(11) =(15/4)*11 + 77/4
Revenue (2008) = 165/4 + 77/4 = 242/4 = $60.50
=========
Cheers,
Stan H.